OpenStudy (boldgeorge):

Medal for who can help! Admission to a dance is $7.00 for a ticket bought in advance and $9.50 for a ticket bought at the door. 300 tickets were sold, for a total of $2405.00. How many tickets of each type were sold? Thanks in advance!

OpenStudy (luigi0210):

set up a system of equations: 7x+9.5y=2405 x+y=300

OpenStudy (luigi0210):

solve for x: x=300-y ------------------------ Sub: 7(300-y)+9.5y=2405 ------------------------- Solve: 2100-7y+9.5y=2405 2.5y=305 y=122 ------------------------ plug back in and solve for x: x+122=300 x=178

OpenStudy (luigi0210):

x=tickets bought in advanced y=tickets bought at door